# EP 250108a/SN 2025kg: Observations of the most nearby Broad-Line Type Ic Supernova Following an Einstein Probe Fast X-ray Transient

### v3

@@ 1 Introduction
Though first discovered in the era of sounding rockets (e.g. Cooke, 1976), and highlighted by discoveries with Chandra and XMM-Newton (Jonker et al., 2013; Alp & Larsson, 2020; Quirola-Vásquez et al., 2023), the nature of fast X-ray transients (“FXTs”; non-repeating transients detected in the X-rays) and their connection to the deaths of massive stars and luminous, jetted -ray bursts (GRBs) remains uncertain. By definition, FXTs represent a broad category of transients. In some cases, FXTs could represent the tail of the emission from classical GRBs whose spectra peak at tens to hundreds of keV. FXTs also encompass the population of so-called X-ray Flashes (Heise et al., 2001; Sakamoto et al., 2005), a historical “sub-type” of GRBs defined by a greater ratio of fluence in the X-ray compared to the -ray bands. Most notably, this XRF/FXT group includes the nearby, serendipitous discoveries of Type Ib/c supernovae (SNe), SN 2006aj (Campana et al., 2006; Pian et al., 2006; Soderberg et al., 2006) identified in the long GRB/XRF 060218 and the X-ray-only discovery of the much less luminous SN 2008D (Soderberg et al., 2008). These discoveries showed that at least some FXTs are produced by the core-collapse of massive, stripped-envelope stars. Though sharing similarities with the broad-line Type Ic (Ic-BL) SNe accompanying GRBs (GRB-SNe), a subset of the SNe following these XRFs were times less luminous compared to the prototypical GRB-SN, GRB 980425/SN 1998bw, and one showed absorption due to He (Soderberg et al., 2006; Modjaz et al., 2009), potentially representing a stage in the continuum of massive star explosions (e.g., Galama et al. 1998; Clocchiatti et al. 2011; Soderberg et al. 2008; Alp & Larsson 2020). However, with a limited sample of such events to date, the mapping of GRBs to FXTs, XRFs or core-collapse SNe without high-energy emission continues to be an open question.
@@
The 2024 launch of a wide-field soft X-ray monitor onboard the Einstein Probe (EP), opened a new route to discover FXTs and characterize their multi-wavelength counterparts (Yuan et al., 2015, 2022). Critically, EP is able to detect FXTs with its Wide Field X-ray Telescope (WXT; 3600 sq. degree field of view), localize them via WXT to arcminute precision, and subsequently re-point a narrow-field follow-up telescope to obtain arcsecond position. As such, the latency of FXT announcements has reduced from days (and, often, year timescales) to minutes (e.g., Zhang et al. 2024; Li et al. 2025). This speed increase offers the unprecedented ability to locate multi-wavelength counterparts and unveil their origins.
@@
On 8 January 2025 at 12:30:28.34 UT, EP discovered a new FXT, EP 250108a, with a duration of s (Li et al., 2025). Later analysis of Fermi-GBM observations place a conservative, sky-averaged upper limit on associated emission in the 10-1000 keV band of erg cm-2 s-1 at 415 s following the initial trigger (Fermi’s view of the localizaton was occulted by Earth prior to this time; Ravasio et al. 2025). Prompt follow-up observations were not taken, but imaging obtained days after the outburst revealed a blue and rapidly-fading counterpart with a redshift of (Eyles-Ferris et al., 2025). Beginning at days following the initial trigger, the cooling, optical counterpart began to rise in brightness. Spectra taken by the Nordic Optical Telescope (NOT) and Gemini-North at days (where refers to the observer-frame time since the EP trigger) showed the signatures of a Ic-BL SN, dubbed SN 2025kg (see Section 2.2 for further analysis; Levan et al. 2025b; Xu et al. 2025). SN 2025kg offers a unique first opportunity for a detailed comparison between the multi-band light curves and spectroscopic properties of FXT SNe, GRB-SNe and stripped-envelope SNe observed without high-energy counterparts, shedding light on FXT progenitors.
@@ 3.2 SN 2025kg Spectra & Broadened H at days
In Figure 5 we show the spectra for GRB-SNe and FXT SNe within a week of the SN peak ( days) and the spectrum of SN 2020bvc taken 1.9 days after the explosion (no later spectra are available on WISEReP). Spectra of the Type Ib SN 2008D show prominent He I 4471, 5876, 6678, 7061 Å absorption lines (Modjaz et al., 2009), which we do not observe in SN 2025kg. Considering the remaining spectra, at both epochs, SN 2010bh provides a poorer fit to SN 2025kg than the SN 2006aj and SN 1998bw, which we attribute in part to its larger ejecta velocities (see Section 3.3; Chornock et al. 2010). Similar to their light curve properties (Section 3.1) SN 1998bw and SN 2006aj are spectroscopically comparable objects to SN 2025kg (Section 2.2). SN 2020bvc also shows a comparable shape to SN 2025kg, in keeping with comparisons between the two objects in our companion paper (Eyles-Ferris et al., 2025)
Turning to the broadened H emission, we compare our SN 2025kg spectrum at days to those of SN 1998bw, SN 2010bh and AT 2018cow at days in the region of H (Figure 6). Spectra of SN 2006aj beyond 20 days of the initial high-energy trigger are not available on WISeREP. Compared to the spectrum of SN 2025kg binned at a resolution of 40 Å, the broadened emission feature is stronger and more narrow in AT 2018cow and absent or very weak in SN 2010bh. Though a comparable bump in the spectrum of SN 1998bw is observed, we note that this likely due to two absorption features on either side of H. Further, in SN 1998bw the spectral shape in this region is is broadly consistent across several weeks of observations. In contrast, in SN 2025kg the continuum is fairly flat over Åat most epochs, but exhibits broad H at one single epoch (Figure 3). Thus, we conclude that absorption features cannot explain the broadened emission feature in SN 2025kg, and that its most likely interpretation is that it arises from H interaction.
@@ 3.4 Near-IR Evidence for Helium
Turning to the JWST spectrum, we observe a prominent, broadened absorption feature 1 m and a shallower absorption feature 2 m (Figure 8). We first explore if the 2 m feature can be explained with only C I m and Mg II m or a combination of the two (i.e., no He I), using a Bayesian toolkit to model blended absorption features from multiple species with Gaussian components (Liu et al., 2023)333https://github.com/slowdivePTG/BayeSpecFit. find velocities in the range of 24,000 - 41,000 km s-1, 10,000 km s-1 faster than the Fe II and Si expansion velocities at a similar epoch (Figure 7). This leads us to disfavor that this feature can be explained by C I and Mg II alone. A two-Gaussian fit of He I and Mg II provide somewhat high (relative to those of Fe II and Si) velocities of 20,000 - 22,000 km s-1, and favors a larger contribution from Mg II than He I. Overall, we conclude that the 2 m feature is likely to be at least partially due to He I absorption, and do not rule out marginal blending with Mg II.
We next consider whether the 1 m feature is due to He I absorption. If both this feature and the 2 m feature are dominated by He I, the expansion and dispersion velocities of both features would be consistent and the equivalent width ratio, EW(He I m)/EW(He I m) (Lucy, 1991). To test these, we first assume this feature is due to He I only, and derive an expansion velocity of 19,000 km s-1, consistent with modeling of the light curve (Section 4) a dispersion velocity of 9,000 km s-1, and a pseudo-EW ratio of 1.47, within the expected range. The larger dispersion velocity of the the 1 m feature relative to the 2 m feature (Figure 9) suggests that the 1 m feature is likely a blend of He I, C I and/or Mg II.
Finally, we simultaneously fit the 1 and 2 m features assuming both could be a combination of the He I lines m and the C I lines m (Figure 9). This fit finds He dominates the 2 m feature and is blended with C I in the 1 m feature. This fit describes the data well and gives a pseudo-EW ratio between the two best-fit He I absorption lines. We conclude that there is likely a small amount of He in the ejecta of SN 2025kg, though a precise estimate is outside the scope of the current work. Blending of He I and another feature is supported by a higher-resolution near-IR spectrum of SN 1998bw spectrum showing two “sub-features” within the 1 m feature (Figure 9). Based on the absence of He I lines in the optical (Figures 2 and 5) we place a conservative upper limit on He in the ejecta of (see, e.g., models for SN Ib/c optical spectra in Dessart et al., 2020).
@@ 3.6 Comparison of SN and High-Energy Properties
In addition to EP 250108a/SN 2025kg, we include the events SNe 1998bw, 2006aj, 2008D, 2010bh, and 2024gsa in this comparison. We utilize the prompt X-ray (2-26 keV) emission of SN 1998bw detected by BeppoSAX (Galama et al., 1998; Pian et al., 2000), X-ray detections by Swift-XRT of SN 2006aj, SN 2008D, and SN 2010bh pulled from UKSSDC (Evans et al., 2007, 2009), and the prompt X-ray emission of SN 2024gsa and SN 2025kg observed by EP (Li et al., 2025; Sun et al., 2024). We caution that these properties are drawn from a variety of instruments, each with different energy ranges and slewing strategies. However, we note that these comparisons span several orders of magnitude, mitigating some biases. We mark the different instruments in Figure 10. For SN properties, we incorporate SN peak magnitudes derived from the light curves presented in Sections 3.1 and Si velocities presented in Section 3.3.
Figure 10 highlights that SN 2025kg, SN 1998bw and SN 2006aj occupy a similar region in the parameter space of peak SN and peak X-ray luminosity (see also Figure 13 of Eyles-Ferris et al. (2025)). We observe a potential trend in these properties for the objects outside of this cluster, though a significantly larger sample is needed to test this trend further. SN 2025kg and SN 2006aj are also comparable in their observed FXT duration. However, we caution that the duration of EP events, in particular, is often set by instrumental rather than astrophysical effects and the duration of EP 250108a is a preliminary value (Li et al., 2025). Of the Swift events, there is a potential trend between SN peak brightness and observed FXT duration. Notably, EP 250108a/SN 2025kg has the longest X-ray duration of the events considered here.
@@ 4.1 One-zone Model
Our model provides a strong fit to the data. We infer of ejecta with a 56Ni mass of and an initial ejecta velocity of km s-1. Our estimate of ejecta mass and energy is consistent with median values inferred from Type Ic-BL SNe (e.g., Taddia et al. 2019; Rodríguez et al. 2023; Srinivasaragavan et al. 2024) using a similar model. However, our estimate of the 56Ni is marginally inconsistent with the Type Ic-BL SNe medial values of Taddia et al. (2019) (at ), which could be due to differences in assumptions about -ray opacities or hinting towards a limitation with the one-zone model.
4.2 Nickel Mixing Model
There are clear limitations with one-zone model in its assumption of a centrally-located 56Ni and constant opacity. In reality, the opacity is almost certainly not gray or constant with time (e.g., Niblett et al., 2025) and 56Ni is likely mixed out to large radii due to Rayleigh-Taylor or jet-driven instabilities (e.g., Khatami & Kasen, 2019; Reichert et al., 2023). The latter effect significantly impacts the overall bolometric light curve. In particular, the classical “Arnett-like” one-zone model often over-predicting the quantity of 56Ni compared to models with mixing (e.g., Bersten et al., 2014) for two reasons. First, the central heating term is forced to overcome adiabatic losses compared to heating from ejecta at larger radii. Second, energy through heating has to diffuse through more ejecta. Therefore, our 56Ni mass measurement above is likely an overestimate.
where is the effective opacity, which smoothly transitions from the maximum opacity, to the minimum opacity, as the temperature starts to approach . This functional form and are chosen to replicate the opacity evolution in detailed numerical simulations (e.g., Nagy, 2018).
We fit this model to the observed light curve as we did in Section 4.1. Our fits with this model are shown in Figure 1 highlighting broad agreement with the data. Compared to the one-zone model, our estimated parameters better match intuition and expectations. In particular, we infer a total ejecta mass of with a 56Ni mass of . We also find evidence for significant mixing, with , that is, 56Ni is distributed out into of the mass layers. The lower 56Ni estimate compared to our one-zone model is driven by the high mixing, which reduces the need for a larger central 56Ni as the mixed material has to overcome less adiabatic losses and diffuse through less ejecta.
Though this model offers several advantages compared to the one described in Section 4, there are still uncertainties, such as the assumed mass profile, therefore, we conservatively use the 56Ni estimate from both models. These masses and mixing provide some strong hints into the progenitor, as we discuss in Section 5. Given the limits of both models, we adopt a conservative range based on both for the 56Ni mass of
@@ 5.1 Constraints on the ZAMS Mass of the Progenitor
Here, we determine the ZAMS mass of the progenitor by combining the inferred masses on the ejecta yield, the 56Ni yield (which dictates the disk mass) and the compact remnant mass. To do this, we set the mass of the star at explosion to the C/O core mass as SN 2025kg shows evidence for stripping of its H and most of its He layers; Section 2.2. We assume that the C/O core mass will be equivalent to the summation of the ejecta mass, the disk mass and the compact object remnant mass. For the first component, we use the results from our light curve fits, which indicate a total ejecta mass of with a 56Ni mass of in SN 2025kg (Section 4). We explore two scenarios to determine the latter two components: a black hole accretion disk jet model and a model invoking a neutron star accretion disk. If the remnant is a black hole, we assume the accretion disk forms around a black hole. If it is a neutron star, we assume an initial remnant mass of .

### v4

@@ 1 Introduction
Though first discovered in the era of sounding rockets (e.g. Cooke, 1976), and highlighted by discoveries with Chandra and XMM-Newton (Jonker et al., 2013; Alp & Larsson, 2020; Quirola-Vásquez et al., 2023), the nature of fast X-ray transients (“FXTs”; non-repeating transients detected in the X-ray) and their connection to the deaths of massive stars and luminous, jetted -ray bursts (GRBs) remains uncertain. By definition, FXTs represent a broad category of transients. In some cases, FXTs could represent the tail of the emission from classical GRBs whose spectra peak at tens to hundreds of keV. FXTs also encompass the population of so-called X-ray Flashes (“XRFs”; Heise et al. 2001; Sakamoto et al. 2005), a historical “sub-type” of GRBs defined by a greater ratio of fluence in the X-ray compared to the -ray bands. Most notably, this XRF/FXT group includes the nearby, serendipitous discoveries of Type Ib/c supernovae (SNe), SN 2006aj (Campana et al., 2006; Pian et al., 2006; Soderberg et al., 2006) identified in the long GRB/XRF 060218 and the X-ray-only discovery of the much less luminous SN 2008D (Soderberg et al., 2008). These discoveries showed that at least some FXTs are produced by the core-collapse of massive, stripped-envelope stars. Though sharing similarities with the broad-line Type Ic (Ic-BL) SNe accompanying GRBs (GRB-SNe), the SNe associated with XRFs demonstrate considerable diversity compared to GRB SNe. For example, two SNe (SN 2008D and SN 2010bh; Soderberg et al. 2008) following these XRFs were times less luminous compared to the prototypical GRB-SN, GRB 980425/SN 1998bw. Further, one (SN 2008D) showed absorption due to He (Soderberg et al., 2006; Modjaz et al., 2009), potentially representing a stage in the continuum of massive star explosions (e.g., Galama et al. 1998; Clocchiatti et al. 2011; Soderberg et al. 2008; Alp & Larsson 2020). However, with a limited sample of such events to date, the mapping of GRBs to FXTs, XRFs or core-collapse SNe without high-energy emission continues to be an open question.
@@
The 2024 launch of a wide-field soft X-ray monitor onboard the Einstein Probe (EP), opened a new route to discover FXTs and characterize their multi-wavelength counterparts (Yuan et al., 2015, 2022). Critically, EP is able to detect FXTs with its Wide Field X-ray Telescope (WXT; 3600 sq. degree field of view), localize them via WXT to arcminute precision, and subsequently re-point a narrow-field follow-up telescope to obtain arcsecond position. As such, the latency of FXT announcements has reduced from days (and, often, year timescales) to minutes (e.g., Zhang et al. 2024; Li et al. 2025a). This speed increase offers the unprecedented ability to locate multi-wavelength counterparts and unveil their origins.
@@
On 8 January 2025 at 12:30:28.34 UT, EP discovered a new FXT, EP 250108a, with a duration of s (Li et al., 2025a, b). Later analysis of Fermi-GBM observations place a conservative, sky-averaged upper limit on associated emission in the 10-1000 keV band of erg cm-2 s-1 at 415 s following the initial trigger (Fermi’s view of the localizaton was occulted by Earth prior to this time; Ravasio et al. 2025). Prompt follow-up observations were not taken, but imaging obtained days after the outburst revealed a blue and rapidly-fading counterpart with a redshift of (Eyles-Ferris et al., 2025). Beginning at days following the initial trigger, the cooling, optical counterpart began to rise in brightness. Spectra taken by the Nordic Optical Telescope (NOT) and Gemini-North at days (where refers to the observer-frame time since the EP trigger) showed the signatures of a Ic-BL SN, dubbed SN 2025kg (see Section 2.2 for further analysis; Levan et al. 2025b; Xu et al. 2025; Srinivasaragavan et al. 2025). SN 2025kg offers a unique first opportunity for a detailed comparison between the multi-band light curves and spectroscopic properties of FXT SNe, GRB-SNe and stripped-envelope SNe observed without high-energy counterparts, shedding light on FXT progenitors.
@@ 3.2 SN 2025kg Spectra & Broadened H at days
In Figure 5 we show the spectra for GRB-SNe and FXT SNe within a week of the SN peak ( days) and the spectrum of SN 2020bvc taken 1.9 days after the explosion (no later spectra are available on WISEReP). Spectra of the Type Ib SN 2008D show prominent He I 4471, 5876, 6678, 7061 Å absorption lines (Modjaz et al., 2009), which we do not observe in SN 2025kg. Considering the remaining spectra, at both epochs, SN 2010bh is less visually comparable to SN 2025kg than SN 2006aj and SN 1998bw, which we attribute in part to its larger ejecta velocities (see Section 3.3; Chornock et al. 2010). Similar to their light curve properties (Section 3.1) SN 1998bw and SN 2006aj are spectroscopically comparable objects to SN 2025kg (Section 2.2). SN 2020bvc also shows a comparable shape to SN 2025kg. We note that SN 2020bvc was also comparable to SN 2025kg in terms of photometric color, blackbody radius, and X-ray properties during the early ( days) fast-cooling phase, as shown in our companion paper (Eyles-Ferris et al., 2025).
Turning to the broadened H emission, we fit a simple Gaussian profile to the feature (omitting the narrow component using curvefit, finding a full width at half maximum value of 50 Å. We next compare our SN 2025kg spectrum at days to those of SN 1998bw, SN 2010bh and AT 2018cow at days in the region of H (Figure 6). Spectra of SN 2006aj beyond 20 days of the initial high-energy trigger are not available on WISeREP. Compared to the spectrum of SN 2025kg binned at a resolution of 40 Å, the broadened emission feature is stronger and more narrow in AT 2018cow and absent or very weak in SN 2010bh. Though a comparable bump in the spectrum of SN 1998bw is observed, we note that this likely due to two absorption features on either side of H. Further, in SN 1998bw the spectral shape in this region is is broadly consistent across several weeks of observations. In contrast, in SN 2025kg the continuum is fairly flat over Åat most epochs, but exhibits broad H at one single epoch (Figure 3). Thus, we conclude that absorption features cannot explain the broadened emission feature in SN 2025kg, and that its most likely interpretation is that it arises from H interaction.
@@ 3.4 Near-IR Evidence for Helium
Turning to the JWST spectrum, we observe a prominent, broadened absorption feature 1 m and, potentially, a shallower absorption feature 2 m (Figure 8). To investigate the source and significance of these features, we utilize a Bayesian toolkit to jointly model the 1 and 2 m features as blended absorption features from multiple species with Gaussian components (Liu et al., 2023)333https://github.com/slowdivePTG/BayeSpecFit. We give the priors for our fits in Appendix Section 6.2. We note that the errors on velocity found through this method do not encompass uncertainties in the fit to the continuum. Thus, the true uncertainties are likely larger than those reported here.
First, we test the significance of the possible feature at 2 m as, if real, the feature would provide a second observational constraint on the presence of He I. We apply two models and compare their and their Bayesian Information Criterion (BIC). The first model fits the 2 m region with a linear function, a reasonable approximation for the continuum in the limited wavelength regime, while the second applies a Gaussian to represent an absorption feature. We find that the Gaussian model results in an improvement in and BIC compared to the linear fit, suggesting that the 2 m feature is real.
We next evaluate fits to the 1 and 2 features for He I, C I and/or Mg II. The larger width of the 1 m feature relative to the 2 m feature (Figure 9) suggests a larger dispersion velocity. In turn, this implies that the 1 m feature is not due to one element alone and is instead likely a blend of He I, C I and/or Mg II. Thus, we first fit both features for He I m and C I m (Figure 9). This fit finds He dominates the 2 m feature and is blended with C I in the 1 m feature. Both the C I (17,600 200 km s-1) and He I (14,500 100 km s-1) expansion velocities are similar to the Si velocities at a similar epoch (Figure 7). A second fit replacing C I with Mg II results in a similar He I velocity and a higher Mg II velocity (21,700 100 km s-1). Finally, we explore if the 1 and 2 m features can be explained with only C I and Mg II or a combination of the two (i.e., no He I). We find expansion velocities in the range of 24,000 - 41,000 km s-1, 10,000 km s-1 faster than the Si expansion velocities at a similar epoch (Figure 7). This finding implies that the 1 and 2 m features are most reasonably fit with the presence of He I.
The most consistent expansion velocities with optical lines (Section 3.3) are derived from fits to these regions that include some He I. We also note that the best-fit pseudo-equivalent width ratio EW(He I m)/EW(He I m) from our first fit is , in line with predictions for these components (Lucy, 1991), also consistent with the presence of He. Blending of He I and another feature is supported by a higher-resolution near-IR spectrum of SN 1998bw spectrum showing two “sub-features” within the 1 m feature (Figure 9).
We conclude that there is likely a small amount of He in the ejecta of SN 2025kg, though a precise estimate is outside the scope of the current work. Based on the absence of He I lines in the optical (Figures 2 and 5) we place a conservative upper limit on He in the ejecta of (see, e.g., models for SN Ib/c optical spectra in Dessart et al., 2020, which are independent of engine type). The extent of He mixing in the ejecta may significantly impact the observed signature (e.g., Dessart et al. 2020, 2024).
@@ 3.6 Comparison of SN and High-Energy Properties
In addition to EP 250108a/SN 2025kg, we include the events SNe 1998bw, 2006aj, 2008D, 2010bh, and 2024gsa in this comparison. We utilize the prompt X-ray (2-26 keV) emission of SN 1998bw detected by BeppoSAX (Galama et al., 1998; Pian et al., 2000), X-ray detections by Swift-XRT of SN 2006aj, SN 2008D, and SN 2010bh pulled from UKSSDC (Evans et al., 2007, 2009), and the prompt X-ray emission of SN 2024gsa and SN 2025kg observed by EP (Sun et al., 2024; Li et al., 2025b). We caution that these properties are drawn from a variety of instruments, each with different energy ranges and slewing strategies. However, we note that these comparisons span several orders of magnitude, mitigating some biases. We mark the different instruments in Figure 10. For SN properties, we incorporate SN peak magnitudes derived from the light curves presented in Sections 3.1 and Si velocities presented in Section 3.3.
Figure 10 highlights that SN 2025kg, SN 1998bw and SN 2006aj occupy a similar region in the parameter space of peak SN and peak X-ray luminosity (see also Figure 13 of Eyles-Ferris et al. (2025)). We observe a potential trend in these properties for the objects outside of this cluster, though a significantly larger sample is needed to test this trend further. In terms of duration, the Swift events show a potential trend between SN peak brightness and observed FXT duration. We do not observe any correlations between the remaining events.
@@ 4.1 One-zone Model
Our model provides a strong fit to the data. Our fits with this model are shown in Figure 1 highlighting broad agreement with the data. We infer of ejecta with a 56Ni mass of and an initial ejecta velocity of km s-1. All uncertainties at credible interval for the one-dimensional marginalized posterior. Our estimate of ejecta mass and energy is consistent with median values inferred from Type Ic-BL SNe (e.g., Taddia et al. 2019; Rodríguez et al. 2023; Srinivasaragavan et al. 2024) using a similar model. However, our estimate of the 56Ni is marginally inconsistent with the Type Ic-BL SNe medial values of Taddia et al. (2019) (at ), which could be due to differences in assumptions about -ray opacities or hinting towards a limitation with the one-zone model. We note that our analysis here is on the photometry, while past works have directly fit a reconstructed bolometric luminosity which could also explain the discrepancy.
4.2 Caveats to One-zone Model
There are clear limitations with one-zone model in its assumption of a centrally-located 56Ni and constant opacity. In reality, the opacity is almost certainly not gray or constant with time (e.g., Niblett et al., 2025) and 56Ni is likely mixed out to large radii due to Rayleigh-Taylor or jet-driven instabilities (e.g., Khatami & Kasen, 2019; Reichert et al., 2023). The latter effect significantly impacts the overall bolometric light curve. In particular, the classical “Arnett-like” one-zone model often over-predicting the quantity of 56Ni compared to models with mixing (e.g., Bersten et al., 2014) for two reasons. First, the central heating term is forced to overcome adiabatic losses compared to heating from ejecta at larger radii. Second, energy through heating has to diffuse through more ejecta. Therefore, our 56Ni mass measurement above is likely an overestimate. Additional luminosity in the form of a central engine could also provide an alternative path to powering the light curve which would naturally reduce the quantity of 56Ni.
In Appendix Section 6.1 we present an alternate model in which 56Ni is mixed with the outer SN layers, resulting in a lower best-fit value for the 56Ni mass (0.2 ). Though this model offers several advantages compared to the one described in Section 4, there are still uncertainties, such as the assumed mass profile, therefore, we conservatively use the 56Ni estimate from both models. These masses and mixing provide some strong hints into the progenitor, as we discuss in Section 5.
Finally, we have assumed that the peak emission is powered entirely by 56Ni decay. Shock heating can contribute to the peak emission, reducing the requirements on the 56Ni mass (Niblett et al., 2025).
@@ 5.1 Constraints on the ZAMS Mass of the Progenitor
Here, we determine the ZAMS mass of the progenitor by combining the inferred masses on the ejecta yield, the 56Ni yield (which dictates the disk mass) and the compact remnant mass. We note that the 56Ni mass may be overestimated due to luminosity contribution from additional sources, but do not anticipate this significantly impacts our order-of-magnitude estimates. To do this, we set the mass of the star at explosion to the C/O core mass as SN 2025kg shows evidence for stripping of its H and most of its He layers; Section 2.2. We assume that the C/O core mass will be equivalent to the summation of the ejecta mass, the disk mass and the compact object remnant mass. For the first component, we use the results from our light curve fits, which indicate a total ejecta mass of with a 56Ni mass of in SN 2025kg (Section 4). We explore two scenarios to determine the latter two components: a black hole accretion disk jet model and a model invoking a neutron star accretion disk. If the remnant is a black hole, we assume the accretion disk forms around a black hole. If it is a neutron star, we assume an initial remnant mass of .